OpenStudy (lizzie.tttt):

What is the slope intercept form of the equation; -3x+4y=8? A; y=3x+2 B; y=3x+8 C;y=3/4x+2 D; y=-3/4x+2

OpenStudy (hippoharry):

−3x+4y=8 Step 1: Add -4y to both sides. −3x+4y+−4y=8+−4y −3x=−4y+8 Step 2: Divide both sides by -3. −3x−3=−4y+8−3x=4 3 y+ −8 3 Answer: x= 4 3 y+ −8 3